LS1 report: According to schedule

The AD’s BHN06 coils have been delivered back to CERN at long last from their repair at a Russian facility. The magnet will be reinstalled underneath the ATRAP experimental area, which has already been partially dismantled to make way for it.

Consolidation of the 7 PS main magnets continues as planned, with 4 magnets already removed from the beam line and delivered to the magnet workshop. A specialised team will be arriving in mid-November (also from Russia) to consolidate the magnets on the surface.

The PS Booster’s beam dump replacement project remains ongoing, with the final survey of the beam transfer line currently underway. The SPS’s irradiated cable replacement campaign continues as planned; it will be completed by the end of March 2014.

Over at the LHC, the Radiation to Electronics (R2E) campaign is progressing well. These activities are actually a few weeks ahead of schedule at Point 1, where teams have already begun commissioning relocated cryogenic equipment.

The LHC’s QRL X-ray campaign continues to search out damaged compensators in sector 2-3. Meanwhile, a damaged compensator has been replaced in sector 1-2, a replacement in sector 5-6 is currently being tested, and a damaged compensator in sector 8-1 has also been replaced and tested successfully.

In sector 4-5, damaged bellows were also found in an electrical feed box. Of the four damaged bellows discovered, one has already been repaired and another will be fixed on the surface during the next weeks.

The Superconducting Magnets and Circuits Consolidation (SMACC) project is advancing. In fact, the first wagon of the SMACC train arrived in sector 4-5 this week. With the opening of the W bellows there, this wagon will have done the full tour of the LHC. This wagon will be reincorporated back into the consolidation train to assist with other SMACC operations.

Elsewhere in the LHC: the first short circuit tests have been carried out in LSS4; electrical maintenance in sector 2-3 has been completed; a non conformity in sector 5/6 is currently being resolved; and a second injector kicker magnet has been installed in LSS8.